### v1

@@ III-F LLM Deployment in Telecom Networks
Practical deployment is the prerequisite for advanced applications of LLM technologies in telecom networks. In particular, it indicates how LLM models can be deployed within the current telecom network architecture, e.g., central cloud, network edge or even user devices. The LLM has great demands for computational and storage resources. For instance, GPT-4 has 1.76 trillion parameters and the model size is 45 GB[63], posing a heavy burden on network storage capacities. Fine-tuning an LLM model with 7 billion parameters, such as GPT-4-LLM[94], could take nearly 3 hours on an 8$\times$80GB A100 machine, which is extremely time-consuming [95]. In addition, the inference time of LLM models will also contribute to overall network latency, which is related to hardware support, batch size, parallelism, model pruning, etc[96]. Therefore, it is of great importance to deploy LLM models appropriately to better serve the telecom network demand. We summarize the existing deployment schemes in Table III, including cloud, network edge, on-device, cache-based, and cooperative deployment. We present the details of each strategy as the following:
@@ III-F1 Cloud deployment
: Considering LLM’s high demand for computational and storage resources, deploying LLM models in the central cloud is a straightforward solution, which can provide substantial computational resources to support the fine-tuning and inference of LLM models[15]. Shen et al. investigate LLM-enabled autonomous edge AI[15], in which the network edge devices can send the user request and datasets feature to LLM models in the cloud, and then the LLM can send back the task planning and AI model configuration to network edge devices through the backhaul. After that, the network edge and user devices can collaborate to make edge inferences. Cloud deployment can easily adapt to existing telecom network architecture, and a few pieces of extra hardware are needed since the LLM is deployed in the virtual cloud. However, cloud deployment suffers from long response time and high bandwidth costs since all data has to be transmitted to the cloud, and then LLM will process the request and finally download the LLM’s output[16]. The long response time may prevent the applications on latency-critical tasks, e.g., vehicle-to-vehicle networks and unmanned aerial vehicle control. In addition, the frequent multimodal information exchange, such as images and videos, between end users and cloud LLM will lead to extra bandwidth costs.
@@ III-F2 Network edge deployment:
Here network edge refers to edge cloud or BSs that are closer to users than central cloud. Network edge deployment can be an appealing approach to shorten the response time and save bandwidth. However, compared with the central cloud, network edge devices usually have limited computational and storage capacities. To this end, multiple techniques can be exploited. For the storage capacity challenge, parameter sharing and model compression may be applied. In particular, LLM models for different downstream tasks may share the same parameters, which can be exploited to save the storage capacity. On the other hand, other technologies may be applied to reduce the computational resources demand in fine-tuning and inference, including parameter-efficient fine-tuning[88], split edge learning[89], and quantized training[90]. With these techniques, deploying LLM models at the network edge becomes a practical strategy.
@@ III-F3 On-device deployment:
There are multiple benefits of deploying LLM models on user-side mobile devices, e.g., fast responses and local customization based on the user’s specific requirements. However, such a deployment is also challenging since LLM models are usually storage- and computation-intensive. Xu et al. introduced a split learning approach based on collaborative end-edge-cloud computing, aiming to deploy LLM agents at mobile devices and network edge [17]. Specifically, the authors assume that LLM models with less than 10B parameters such as LLAMA-7B can operate on mobile devices, providing real-time inference services. Meanwhile, LLM models with more than 10B parameters such as GPT-4 are deployed on network edge servers, using global information and historical memory to assist the mobile LLM in processing complex tasks. Such a collaboration enables higher flexibility by exploiting mobile LLM models. However, the study of on-device LLM is still in a very early stage, and it requires considerable efforts to prove the feasibility of such a design[97]. For instance, Apple has proposed a technique to store LLM model parameters on flash memory[91], achieving a 20 times faster inference speed than using GPU with limited dynamic random-access memory (DRAM) capacity. Similarly, Qualcomm has recently announced the Snapdragon 8s Gen 3 mobile platform, which supports popular small-scale LLM such as Llama 2 and Gemini Nano[92]. These studies may pave the way to effective inference of on-device LLMs.
@@ III-F5 Cooperative deployment
Finally, Fig. 3 illustrates different LLM deployment strategies. Note that these approaches may be combined, e.g., deploying small-scale on-device LLMs and then using larger cloud models to update the local models periodically. Given these deployment methods, many critical problems can then be investigated, e.g., service delay evaluation and task offloading, which still require more research efforts. For example, Chen et al. proposed a NETGPT scheme in [98], involving offload architecture, splitting architecture, and synergy architecture for cloud-edge collaboration.
@@ III-G Analyses of LLM Fundamentals in the Telecom Domain
For telecom applications, pre-training an LLM model from scratch can be time-consuming. It first requires extensive dataset collection, and the dataset preprocessing has to consider the format of complicated telecom equations and theories. Meanwhile, it also requires considerable computational resources to pre-train LLM models, leading to heavy burdens for telecom networks. By contrast, a more efficient approach is to fine-tune a general-domain LLM model for specific telecom-domain tasks. Applying LLM technologies to the telecom domain requires an in-depth understanding of these fine-tuning techniques, such as instruction and alignment tuning methods. In particular, instruction tuning involves carefully constructing and selecting instruction datasets, employing strategic tuning methodologies, and considering practical implementation aspects. These strategies will significantly improve the performance, generalization, and user alignment of LLM technologies in the telecom domain. On the other hand, alignment tuning is a multifaceted process involving the setting of ethical guidelines, collection of human feedback, and application of advanced fine-tuning techniques such as RLHF. However, adapting these state-of-the-art fine-tuning techniques to telecom environments is still an open question. The fine-tuning process is usually task-specific, which requires professional knowledge of various telecom domain tasks. Instruction tuning can be a promising method for building a telecom-LLM by using existing telecom knowledge, but the dataset collection can be difficult due to privacy issues.
@@
Prompting techniques are especially useful for solving real-time telecom tasks with stringent delay requirements, e.g., resource allocation and user association. It means that LLM models can directly learn from the inputs and generate desired outputs without extra training, avoiding the tedious model training process in conventional ML algorithms. For instance, ICL provides a framework for leveraging the LLM in new task domains without explicit retraining, with its effectiveness heavily influenced by the design and structure of demonstrations. Meanwhile, CoT prompting has emerged as a potent method for eliciting deeper reasoning capabilities in LLM models, applicable to a range of complex reasoning tasks. While still evolving, this approach opens new avenues for LLM application across diverse problem domains such as the telecom field. In addition, prompt-based planning represents a sophisticated approach to navigating complex tasks, enhancing LLM’s problem-solving capabilities through structured action sequences, feedback integration, and continuous plan refinement. Such planning capabilities are very important for telecom applications since many telecom tasks involve multi-step thinking with complicated procedures. For instance, the resource allocation may include multi-layer controllers[4], and optimization problems can involve several agents and elements[99]. Therefore, multi-step planning and thinking should be carefully designed for LLM-enabled telecom applications.
@@
Evaluation metrics are critical to assess the LLM model’s performance in telecom environments. For instance, efficiency is one of the most important metrics that should be considered in telecom applications since many tasks require rapid or even real-time responses. Therefore, LLM models with long inference times may be inappropriate for these mission-critical applications, e.g., Ultra-Reliable Low Latency Communications (URLLC). In addition, evaluating the performance of LLM models should also include their proneness to hallucination and ethical standards, e.g., LLM may make misleading or even wrong decisions in network management. As LLM design and models continue to evolve and integrate more deeply into various aspects of society, the criteria for their evaluation will likely expand and become more sophisticated. Ensuring that LLM models are accurate, reliable, efficient, and ethically responsible is essential for their sustainable and beneficial integration into human-centric applications.
@@ IV LLM for Generation Problems in Wireless Networks
Domain knowledge generation [100] Adapting a BERT-like model to the telecom domain and testing the model performance by question answering downstream task in the target domain. The proposed technique achieved $F1$ score of 61.20 and $EM$ score of 36.48 on question answering in a small-scale telecom question answering dataset. Techniques such as customizing LLM models to understand and apply telecom-specific language, evaluating their genuine understanding of domain knowledge generation can contribute to more efficient, reliable, and secure telecom service applications. Existing studies have demonstrated the capability of LLM techniques to be applied in telecom, including question answering, literature review, generating troubleshooting report. It shows great promises to build next generation communication networks.
@@
[102] Question answering test on various LLM models, e.g., GPT 3.5, GPT 4, and Bard, including telecom knowledge and product questions. Bard and GPT4 show promise with respect to accuracy and could be useful for telecom domain question and answering. LLM’s summarization requires reliability tests.
[103] Integrating domain-specific grammars into LLM models to guide the generation of structured language outputs, enhancing performance in domain-specific tasks. Demonstrating the efficacy of integrating domain-specific grammars with LLM models in enhancing their ability to generate structured language outputs tailored to specific domains. It emphasizes the potential of this approach to significantly improve LLM performance in domain-specific tasks.
Code generation [14] Using LLM models to generate Verilog code for wireless communication system development in FPGA. The experiment was implemented in the OpenWiFi project. The LLM is capable of refactoring, reusing, and validating existing code. With proper design and prompting, LLM models can generate more complicated projects with multi-step scheduling. LLM greatly reduced the coding time of undergraduate and graduate students by 65.16% and 68.44%, respectively. Code is the cornerstone of modern communication networks, and the LLM provide promising opportunities to improve the efficiency and reliability of codes, and meanwhile greatly save human effort. a) The LLM can refactor and validate existing code. This is very useful in telecom filed, since the network architecture is constantly evolving and updated; b) With proper prompting, the LLM can generate complicated projects with multi-step scheduling requirements, which is very common in telecom filed due to complicated network elements with diverse functions.
@@
[18] Employing four students to reproduce the results of existing network studies with the assistance of LLM models. The students successfully reproduced networking systems by prompting engineering ChatGPT. They also achieve much lower lines of code by using ChatGPT, e.g., one of them is only 20% of the open-source existing version.
@@
[106] Fine-tuning pre-trained LLM models on Verilog datasets collected from GitHub and Verilog textbooks and then generating Verilog projects. Fine-tuning LLM models over a specific language can improve the coding correct rate by 26%.
Network configuration generation [107] It proposed a three-stage LLM-aided progressive policy generation pipeline for intent decomposition. Through evaluating a service chain use case, the paper found LLM models could generalize to new intents through few-shot learning and concluded leveraging LLM models for policy generation is promising for automatic intent-based application management. Telecom network operators can leverage the LLM for network configuration generation in various ways. This includes automatic network provisioning, optimization and performance tuning, security and compliance configuration, fault diagnosis and troubleshooting, and network virtualization. The LLM enables efficient, reliable, and secure generation of network configurations, reducing manual effort and improving network management in telecom environments.
[108] It proposed a multi-stage framework that utilizes LLM models to automate network configuration by taking in natural language requirements and translating them into formal specification, high-level configurations, and low-level device configurations. The results showed that state-of-the-art LLM technologies like GPT-4 are capable of generating fully working configurations from natural language requirements without any fine-tuning.
[109] It proposed a framework that combines LLM models with verifiers, using localized feedback from verifiers to automatically correct errors in configurations generated by the LLM. The proposed scheme is able to synthesize reasonable though imperfect configurations with significantly reduced human effort, and coupling LLM models with verifiers providing localized feedback is necessary for real-world use configurations despite requiring more testing.
@@ IV-A Motivations of Using LLM-based Generation for Telecom
This subsection will introduce the key motivations of using LLM-enabled generation for telecom applications. Firstly, LLM can make telecom knowledge more accessible. LLM models have been pre-trained on many real-world datasets and equipped with considerable knowledge from various fields. Therefore, question-answering has become the most well-known application of LLMs. With domain-specific datasets from websites and textbooks, the LLM can extract professional knowledge from existing publications and then generate appropriate answers based on users’ requests. For instance, Maatouk et al. build a telecom knowledge dataset in [110], including 25,000 pages from research publications, overview, and standards specifications. With proper training and fine-tuning, such a dataset can greatly contribute to a Telecom-GPT, providing a systematic overview of hundreds of publications and standards. With reasoning and comprehension capabilities, professional telecom knowledge will become much more accessible to all researchers and even benefit the whole society.
Meanwhile, LLM’s generation capabilities can also automate many tasks that are usually time-consuming. For instance, developing new standard specifications usually requires considerable writing, discussions, and reviews. By contrast, given enough historical reports and proper prompts, the LLM can produce a draft standard instantly, and then the experts can review it accordingly. Moreover, the experts’ comments can be fed directly to LLM models, and then the LLM can produce a new version efficiently, significantly saving human efforts on writing and revising paper works. Similarly, LLM technologies have been used to generate code in many existing studies, which is one of the most time-consuming tasks of modern industry[18, 105]. LLM models can refactor and improve existing codes, contributing to developing telecom projects.
In addition, LLM models can easily learn from the provided existing examples, which is known as ICL. This capability is particularly useful in generation tasks, and LLM models can quickly generalize the given examples to related unseen scenarios. Meanwhile, if the initial generated output can not satisfy the requirements, users can also send the feedback directly to the LLM input, and then the LLM agent will revise the generation accordingly. This user-friendly generation approach will lower the difficulty of applying LLM techniques to generation tasks in telecom, which usually requires considerable professional knowledge and experience.

### v2

@@ III-F LLM Deployment in Telecom Networks
Practical deployment is the prerequisite for advanced applications of LLM technologies in telecom networks. In particular, it indicates how LLMs can be deployed within the current telecom network architecture, e.g., central cloud, network edge or even user devices. The LLM has great demands for computational and storage resources. For instance, GPT-4 has 1.76 trillion parameters and the model size is 45 GB[73], posing a heavy burden on network storage capacities. Fine-tuning an LLM with 7 billion parameters, such as GPT-4-LLM[109], could take nearly 3 hours on an 8$\times$80GB A100 machine, which is extremely time-consuming [110]. In addition, the inference time of LLMs will also contribute to overall network latency, which is related to hardware support, batch size, parallelism, model pruning, etc[111]. Therefore, it is of great importance to deploy LLMs appropriately to better serve the telecom network demand. We summarize the existing deployment schemes in Table III, including cloud, network edge, on-device, cache-based, and cooperative deployment. We present the details of each strategy as the following:
@@ III-F1 Cloud deployment
: Considering LLM’s high demand for computational and storage resources, deploying LLMs in the central cloud is a straightforward solution, which can provide substantial computational resources to support the fine-tuning and inference of LLMs[15]. Shen et al. investigate LLM-enabled autonomous edge AI[15], in which the network edge devices can send the user request and datasets feature to LLMs in the cloud, and then the LLM can send back the task planning and AI model configuration to network edge devices through the backhaul. After that, the network edge and user devices can collaborate to make edge inferences. Cloud deployment can easily adapt to existing telecom network architecture, and a few pieces of extra hardware are needed since the LLM is deployed in the virtual cloud. However, cloud deployment suffers from long response time and high bandwidth costs since all data has to be transmitted to the cloud, and then LLM will process the request and finally download the LLM’s output[16]. The long response time may prevent the applications on latency-critical tasks, e.g., vehicle-to-vehicle networks and unmanned aerial vehicle control. In addition, the frequent multimodal information exchange, such as images and videos, between end users and cloud LLM will lead to extra bandwidth costs.
@@ III-F2 Network edge deployment:
Here network edge refers to edge cloud or BSs that are closer to users than central cloud. Network edge deployment can be an appealing approach to shorten the response time and save bandwidth. However, compared with the central cloud, network edge devices usually have limited computational and storage capacities. To this end, multiple techniques can be exploited. For the storage capacity challenge, parameter sharing and model compression may be applied. In particular, LLMs for different downstream tasks may share the same parameters, which can be exploited to save the storage capacity. On the other hand, other technologies may be applied to reduce the computational resources demand in fine-tuning and inference, including parameter-efficient fine-tuning[103], split edge learning[104], and quantized training[105]. With these techniques, deploying LLMs at the network edge becomes a practical strategy.
@@ III-F3 On-device deployment:
There are multiple benefits of deploying LLMs on user-side mobile devices, e.g., fast responses and local customization based on the user’s specific requirements. However, such a deployment is also challenging since LLMs are usually storage- and computation-intensive. Xu et al. introduced a split learning approach based on collaborative end-edge-cloud computing, aiming to deploy LLM agents at mobile devices and network edge [17]. Specifically, the authors assume that LLMs with less than 10B parameters such as LLAMA-7B can operate on mobile devices, providing real-time inference services. Meanwhile, LLMs with more than 10B parameters such as GPT-4 are deployed on network edge servers, using global information and historical memory to assist the mobile LLM in processing complex tasks. Such a collaboration enables higher flexibility by exploiting mobile LLMs. However, the study of on-device LLM is still in a very early stage, and it requires considerable efforts to prove the feasibility of such a design[112]. For instance, Apple has proposed a technique to store LLM parameters on flash memory[106], achieving a 20 times faster inference speed than using GPU with limited dynamic random-access memory (DRAM) capacity. Similarly, Qualcomm has recently announced the Snapdragon 8s Gen 3 mobile platform, which supports popular small-scale LLM such as Llama 2 and Gemini Nano[107]. These studies may pave the way to effective inference of on-device LLMs.
@@ III-F5 Cooperative deployment
Finally, Fig. 4 illustrates different LLM deployment strategies. Note that LLM’s requirements for storage and computational resources are the main motivations for developing various deployment strategies. For instance, the model size of Llama3-8b is around 5 GB, and therefore it is possible to be implemented at the network edge or even user devices, i.e., Snapdragon 8s Gen 3 mobile platform recently developed by Qualcomm. Similarly, Gemini Nano is less than 2 GB, and such a small size allows on-device deployment, e.g., Google plans to load Gemini Nano to its Pixel 8 smartphones. By contrast, large-scale LLMs require much more computation resources. For example, inference with Llama3-70b consumes at least 140 GB of GPU RAM. Using 2-bit quantization, the Llama3-70b can be implemented on a 24 GB consumer GPU, but such a low-precision quantization will significantly degrade the model accuracy. To this end, hybrid deployment methods such as cache-based and cooperative deployment are proposed. The key objective is to take advantage of large-scale LLM’s high accuracy, while reducing the dependency on computational resources. On the other hand, these approaches may be combined, e.g., deploying small-scale on-device LLMs and then using larger cloud models to update the local models periodically. Given these deployment methods, many critical problems can then be investigated, e.g., service delay evaluation and task offloading, which still require more research efforts. For example, Chen et al. proposed a NETGPT scheme in [113], involving offload architecture, splitting architecture, and synergy architecture for cloud-edge collaboration.
@@ III-G Analyses of LLM Fundamentals in the Telecom Domain
For telecom applications, pre-training an LLM from scratch can be time-consuming. It first requires extensive dataset collection, and the dataset preprocessing has to consider the format of complicated telecom equations and theories. Meanwhile, it also requires considerable computational resources to pre-train LLMs, leading to heavy burdens for telecom networks. By contrast, a more efficient approach is to fine-tune a general-domain LLM for specific telecom-domain tasks. Applying LLM technologies to the telecom domain requires an in-depth understanding of these fine-tuning techniques, such as instruction and alignment tuning methods. In particular, instruction tuning involves carefully constructing and selecting instruction datasets, employing strategic tuning methodologies, and considering practical implementation aspects. These strategies will significantly improve the performance, generalization, and user alignment of LLM technologies in the telecom domain. On the other hand, alignment tuning is a multifaceted process involving the setting of ethical guidelines, collection of human feedback, and application of advanced fine-tuning techniques such as RLHF. However, adapting these state-of-the-art fine-tuning techniques to telecom environments is still an open question. The fine-tuning process is usually task-specific, which requires professional knowledge of various telecom domain tasks. Instruction tuning can be a promising method for building a telecom-LLM by using existing telecom knowledge, but the dataset collection can be difficult due to privacy issues.
@@
Prompting techniques are especially useful for solving real-time telecom tasks with stringent delay requirements, e.g., resource allocation and user association. It means that LLMs can directly learn from the inputs and generate desired outputs without extra training, avoiding the tedious model training process in conventional ML algorithms. For instance, ICL provides a framework for leveraging the LLM in new task domains without explicit retraining, with its effectiveness heavily influenced by the design and structure of demonstrations. Meanwhile, CoT prompting has emerged as a potent method for eliciting deeper reasoning capabilities in LLMs, applicable to a range of complex reasoning tasks. While still evolving, this approach opens new avenues for LLM application across diverse problem domains such as the telecom field. In addition, prompt-based planning represents a sophisticated approach to navigating complex tasks, enhancing LLM’s problem-solving capabilities through structured action sequences, feedback integration, and continuous plan refinement. Such planning capabilities are very important for telecom applications since many telecom tasks involve multi-step thinking with complicated procedures. For instance, the resource allocation may include multi-layer controllers[4], and optimization problems can involve several agents and elements[114]. Therefore, multi-step planning and thinking should be carefully designed for LLM-enabled telecom applications.
@@
Evaluation metrics are critical to assess the LLM’s performance in telecom environments. For instance, efficiency is one of the most important metrics that should be considered in telecom applications since many tasks require rapid or even real-time responses. Therefore, LLMs with long inference times may be inappropriate for these mission-critical applications, e.g., Ultra-Reliable Low Latency Communications (URLLC). In addition, evaluating the performance of LLMs should also include their proneness to hallucination and ethical standards, e.g., LLM may make misleading or even wrong decisions in network management. As LLM design and models continue to evolve and integrate more deeply into various aspects of society, the criteria for their evaluation will likely expand and become more sophisticated. Ensuring that LLMs are accurate, reliable, efficient, and ethically responsible is essential for their sustainable and beneficial integration into human-centric applications.
@@ IV LLM for Generation Problems in Wireless Networks
Domain knowledge generation [115] Adapting a BERT-like model to the telecom domain and testing the model performance by question answering downstream task in the target domain. The proposed technique achieved $F1$ score of 61.20 and $EM$ score of 36.48 on question answering in a small-scale telecom question answering dataset. Techniques such as customizing LLMs to understand and apply telecom-specific language, evaluating their genuine understanding of domain knowledge generation can contribute to more efficient, reliable, and secure telecom service applications. Existing studies have demonstrated the capability of LLM techniques to be applied in telecom, including question answering, literature review, generating troubleshooting report. It shows great promises to build next generation communication networks.
@@
[117] Question answering test on various LLMs, e.g., GPT 3.5, GPT 4, and Bard, including telecom knowledge and product questions. Bard and GPT4 show promise with respect to accuracy and could be useful for telecom domain question and answering. LLM’s summarization requires reliability tests.
[118] Integrating domain-specific grammars into LLMs to guide the generation of structured language outputs, enhancing performance in domain-specific tasks. Demonstrating the efficacy of integrating domain-specific grammars with LLMs in enhancing their ability to generate structured language outputs tailored to specific domains. It emphasizes the potential of this approach to significantly improve LLM performance in domain-specific tasks.
Code generation [14] Using LLMs to generate Verilog code for wireless communication system development in FPGA. The experiment was implemented in the OpenWiFi project. The LLM is capable of refactoring, reusing, and validating existing code. With proper design and prompting, LLMs can generate more complicated projects with multi-step scheduling. LLM greatly reduced the coding time of undergraduate and graduate students by 65.16% and 68.44%, respectively. Code is the cornerstone of modern communication networks, and the LLM provide promising opportunities to improve the efficiency and reliability of codes, and meanwhile greatly save human effort. a) The LLM can refactor and validate existing code. This is very useful in telecom filed, since the network architecture is constantly evolving and updated; b) With proper prompting, the LLM can generate complicated projects with multi-step scheduling requirements, which is very common in telecom filed due to complicated network elements with diverse functions.
@@
[18] Employing four students to reproduce the results of existing network studies with the assistance of LLMs. The students successfully reproduced networking systems by prompting engineering ChatGPT. They also achieve much lower lines of code by using ChatGPT, e.g., one of them is only 20% of the open-source existing version.
@@
[121] Fine-tuning pre-trained LLMs on Verilog datasets collected from GitHub and Verilog textbooks and then generating Verilog projects. Fine-tuning LLMs over a specific language can improve the coding correct rate by 26%.
Network configuration generation [122] It proposed a three-stage LLM-aided progressive policy generation pipeline for intent decomposition. Through evaluating a service chain use case, the paper found LLMs could generalize to new intents through few-shot learning and concluded leveraging LLMs for policy generation is promising for automatic intent-based application management. Telecom network operators can leverage the LLM for network configuration generation in various ways. This includes automatic network provisioning, optimization and performance tuning, security and compliance configuration, fault diagnosis and troubleshooting, and network virtualization. The LLM enables efficient, reliable, and secure generation of network configurations, reducing manual effort and improving network management in telecom environments.
[123] It proposed a multi-stage framework that utilizes LLMs to automate network configuration by taking in natural language requirements and translating them into formal specification, high-level configurations, and low-level device configurations. The results showed that state-of-the-art LLM technologies like GPT-4 are capable of generating fully working configurations from natural language requirements without any fine-tuning.
[124] It proposed a framework that combines LLMs with verifiers, using localized feedback from verifiers to automatically correct errors in configurations generated by the LLM. The proposed scheme is able to synthesize reasonable though imperfect configurations with significantly reduced human effort, and coupling LLMs with verifiers providing localized feedback is necessary for real-world use configurations despite requiring more testing.
@@ IV-A Motivations of Using LLM-based Generation for Telecom
This subsection will introduce the key motivations of using LLM-enabled generation for telecom applications. Firstly, LLM can make telecom knowledge more accessible. LLMs have been pre-trained on many real-world datasets and equipped with considerable knowledge from various fields. Therefore, question-answering has become the most well-known application of LLMs. With domain-specific datasets from websites and textbooks, the LLM can extract professional knowledge from existing publications and then generate appropriate answers based on users’ requests. For instance, Maatouk et al. build a telecom knowledge dataset in [125], including 25,000 pages from research publications, overview, and standards specifications. With proper training and fine-tuning, such a dataset can greatly contribute to a Telecom-GPT, providing a systematic overview of hundreds of publications and standards. With reasoning and comprehension capabilities, professional telecom knowledge will become much more accessible to all researchers and even benefit the whole society.
Meanwhile, LLM’s generation capabilities can also automate many tasks that are usually time-consuming. For instance, developing new standard specifications usually requires considerable writing, discussions, and reviews. By contrast, given enough historical reports and proper prompts, the LLM can produce a draft standard instantly, and then the experts can review it accordingly. Moreover, the experts’ comments can be fed directly to LLMs, and then the LLM can produce a new version efficiently, significantly saving human efforts on writing and revising paper works. Similarly, LLM technologies have been used to generate code in many existing studies, which is one of the most time-consuming tasks of modern industry[18, 120]. LLMs can refactor and improve existing codes, contributing to developing telecom projects.
In addition, LLMs can easily learn from the provided existing examples, which is known as ICL. This capability is particularly useful in generation tasks, and LLMs can quickly generalize the given examples to related unseen scenarios. Meanwhile, if the initial generated output can not satisfy the requirements, users can also send the feedback directly to the LLM input, and then the LLM agent will revise the generation accordingly. This user-friendly generation approach will lower the difficulty of applying LLM techniques to generation tasks in telecom, which usually requires considerable professional knowledge and experience.